Hello i am having issue with this chapter in ECO4223 if any one can help would be great!

A bond has a face value of $800 and a 10% coupon rate, its current price is $740, and is expected to increase to $750 next year.

The current yield is __% (Enter response rounded to one decimal place.)
Answer #1

Face value = 800

Coupon rate = 10%

Coupon payment = 0.1 * 800 = 80

Sale price this year = 740

Price next year = 750

Total return in one year = (750 - 740) + 80 = 10 + 80 = 90

Current yeild = 90 / 740 * 100 = 12.1621% = 12.2%